OpenStudy (anonymous):

Write and solve the system of equations below. Let a represent adults and let c represent children. 1,200 tickets were sold to the spring carnival. Adult tickets cost $6.00 and children's tickets cost $3.50. A total of $6,450 was collected. How many tickets of each kind were sold? A. a = 800, c = 400 B. a = 900, c = 300 C. a = 1,000, c = 200 D. a = 1,100, c = 100

OpenStudy (anonymous):

ok if u have above conditions .. u can have an equation .. 3.50*c+6*a=6450

OpenStudy (anonymous):

and the other equation is a+c=1200

OpenStudy (anonymous):

so it B ?

OpenStudy (anonymous):

c=1100 ?

OpenStudy (anonymous):

now u have two equations with just two variables .. put the value of variable a=1200-c in 1st equation and get the answer for c by eliminating a from equation

OpenStudy (anonymous):

now put it in above equation and c what comes out for c 3.5*c+6(1200-c)=6450

OpenStudy (anonymous):

now solve it and tell me the answer

OpenStudy (anonymous):

B is answer man

OpenStudy (anonymous):

im a girl dude ! and thats what i asked you .

OpenStudy (anonymous):

can`t u solve it tell me .. i will show u all calculation then

OpenStudy (anonymous):

ok buddy !! u got it ???

OpenStudy (anonymous):

solve at least the last equation u can see c=300 .. ok u should learn in this problem first to make simultaneous equations then solve them by elimination one variable as i did by eliminating a and got c.. then by putting c i got a too..